17May2007 Observer: Peter Buschkamp, Nicolas Bouche TO: David Gonzalez

20:00 we are ready, LBC is ready, the telescope seems ready, but rain is coming our way

21:00 still drops of rain coming down with stars overhead?!?

21:30 looks like we are opening the dome! CB clouds dissipating around us. Encoders reset. MCS crashed.

23:00 still closed. Thick cirrus.

23:30 MCS still crashed. Distant thunder+lightning. Sky clear. Humidity rising quickly 75%.

00:10 MCS fixed. Try focus.

00:30 telescope in focus; seeing 0.88" but clouds.

01:00 threatening clouds passing overhead. seeing 1.4"

01:30 closed

02:30 reopen. Trying on the RX17_1 cluster. seeing 3"->4"

03:00 seeing 1.5" RX17_1 repeat. seeing 1.3"-1.1" (U band)

03:30 RX17_4 2 exp. twilights coming.

04:00 standards + flats.

-- NicolasBouche - 18 May 2007
Topic revision: r4 - 10 Jun 2007, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback